Make a sketch for each problem. Approximately each square root to the nearest hundredth.
Find the area of a triangle with three sides of length 4 cm.

Answer by stanbon(75887)

Make a sketch for each problem. Approximately each square root to the nearest hundredth.
Find the area of a triangle with three sides of length 4 cm.
---------------
Sketch the equilateral triangle.
Sketch an altitude from the top vertex to the base opposite.
That is a perpendicular bisector of the base.
---------
You now have two right triangles with base = 2 and hypotenuse = 4
The altitide = sqrt[4^2-2^2] = sqrt(14)
---------
Area of the original triangle = 2*(1/2)base*height = b*h
= 2*sqrt(14) = 7.48 cm^2
------
Cheers,
Stan H.
-----------------